OpenStudy (anonymous):

Integrate x^2/((x-1) (x+1) (x+1)) dx Can someone please help me?

OpenStudy (anonymous):

partial fractions

OpenStudy (anonymous):

How do I do the plug in method? Where I plug in a value to solve for the variables.

OpenStudy (anonymous):

I got to: A(x+1)^2 +B(x+1)(x-1) + (Cx+D)(x-1)=x^2

OpenStudy (anonymous):

https://www.youtube.com/watch?v=_YMG5mQTAOg This guy is really good at explaining it.

OpenStudy (anonymous):

It's how I learned

OpenStudy (anonymous):

Take the integral: integral x^2/((x-1) (x+1)^2) dx For the integrand x^2/((x-1) (x+1)^2), use partial fractions: = integral (3/(4 (x+1))-1/(2 (x+1)^2)+1/(4 (x-1))) dx Integrate the sum term by term and factor out constants: = 3/4 integral 1/(x+1) dx-1/2 integral 1/(x+1)^2 dx+1/4 integral 1/(x-1) dx For the integrand 1/(x+1), substitute u = x+1 and du = dx: = 3/4 integral 1/u du-1/2 integral 1/(x+1)^2 dx+1/4 integral 1/(x-1) dx The integral of 1/u is log(u): = (3 log(u))/4-1/2 integral 1/(x+1)^2 dx+1/4 integral 1/(x-1) dx For the integrand 1/(x+1)^2, substitute s = x+1 and ds = dx: = (3 log(u))/4-1/2 integral 1/s^2 ds+1/4 integral 1/(x-1) dx The integral of 1/s^2 is -1/s: = 1/(2 s)+(3 log(u))/4+1/4 integral 1/(x-1) dx For the integrand 1/(x-1), substitute p = x-1 and dp = dx: = 1/(2 s)+(3 log(u))/4+1/4 integral 1/p dp The integral of 1/p is log(p): = (log(p))/4+1/(2 s)+(3 log(u))/4+constant Substitute back for p = x-1: = (3 s log(u)+s log(x-1)+2)/(4 s)+constant Substitute back for s = x+1: = (3 (x+1) log(u)+(x+1) log(x-1)+2)/(4 (x+1))+constant Substitute back for u = x+1: = ((x+1) log(x-1)+3 (x+1) log(x+1)+2)/(4 (x+1))+constant Which is equal to: Answer: = 1/4 (2/(x+1)+log(x-1)+3 log(x+1))+constant

OpenStudy (anonymous):

sorry let me clean that up

OpenStudy (anonymous):

Wait.... How did you get the 3/4, 1/2 and 1/4?

OpenStudy (anonymous):

That's my question....

OpenStudy (anonymous):

I think I understand now...

OpenStudy (anonymous):

Thanks guys!!

OpenStudy (anonymous):

Sure thing! good luck!

OpenStudy (anonymous):

sorry i wasn't too helpful lol